OpenStudy (anonymous):

What is the directrix of the parabola with the equation y+3=1/10(x+2)squared?

OpenStudy (jdoe0001):

site is a bit laggy, thus \(\bf 4p(y-{\color{blue}{ k}})=(x-{\color{brown}{ h}})^2\qquad vertex\ ( {\color{brown}{ h}},{\color{blue}{ k}})\\ \quad \\ \quad \\ y+3=\cfrac{1}{10}(x+2)^2\implies y+3=\cfrac{(x+2)^2}{10}\implies 10(y+3)=(x+2)^2\) see the vertex of the parabola?

OpenStudy (anonymous):

No I'm sorry.

OpenStudy (jdoe0001):

\(\bf y+3=\cfrac{1}{10}(x+2)^2\implies y+3=\cfrac{(x+2)^2}{10}\implies 10(y+3)=(x+2)^2 \\ \quad \\ 10(y-(-3))=(x-(-2))^2\) what about now?